end . is the r than, ithin 44. Calculate Water flows with a speed of 1.3 m/s through a section of hose with a cross-sectional area of 0.0075 m2. If the cross-sectional area narrows down to 0.0033 m2 what is the new speed of the water?
Answer #1

here,

the cross sectional area , A0 = 0.0075 m^2

v0 = 1.3 m/s

the new transaltional speed be v

the new cross sectional area , A = 0.0033 m^2

using equation of continuity

A0 * v0 = A * v

0.0075 * 1.3 = 0.0033 * v

solving for v

v = 2.95 m/s

the new speed is 2.95 m/s